OpenStudy (anonymous):

I need some help, am totally lost... Use the given information to evaluate each expression. sin α =√(3 /2) with 90° < α < 180° (a) sin(α/2) (b) cos(α/2) (c) tan(α/2)

OpenStudy (anonymous):

you know \(\alpha =120\) since \(\sin(120)=\frac{\sqrt{3}}{2}\)

OpenStudy (anonymous):

your job is to know find \(\sin(60),\cos(60),\tan(60)\)

OpenStudy (anonymous):

*now is what i meant, although it would also be good to "know" those values

OpenStudy (anonymous):

Thank You so much, it geared me in the right direction! :)

OpenStudy (anonymous):

yw